# Elements of Cohomology for model-theorically finite-dimensional groups and Lie algebras [^1] Keywords: nilpotent group, nilpotent Lie algebra, cohomology, finite-dimensional theory MSC 2020 : 17B56, 03C60

### v1

@@ Abstract
We use the cohomology theory for groups invented by Hochshild and Serre to compute the first cohomology group for nilpotent groups that are definable in a finite-dimensional theory. Based on the established cohomological results, we derive some structural results : we prove a weak form of the Frattini argument for definable connected Cartan subgroups and we give a definable version of Maschke’s theorem in the case of a definable abelian $p^{\perp}$-group. The same results hold when one works with Lie algebras instead of groups.
@@ 1 Introduction
Under the impetus of Hochschild and Serre (see [7] and [8]), techniques from homological algebra have been successfully applied to the study of modules in the category of groups and of Lie algebras. For groups, cohomological methods play a prominent role in the classification of group extensions and had striking applications in number theory through Galois cohomology (see [10]). In the case of Lie algebras, a number of structural results on finite-dimensional (in the usual linear sense) solvable Lie algebras were obtained by Barnes by using cohomological tools (see [2], [3], and [4]).
Therefore, it would be natural to generalize these methods to the context of groups/Lie algebras and associated modules when one works within the category of definable objects in a given first-order theory. We will consider the case of finite-dimensional theories (precise definitions will be given in the following). The existence of a finite dimension on the collection of interpretable sets, endowed in addition with good algebraic properties generalizing those of linear dimension, will enable us to have better control over constructions arising from homological algebra.
Regarding the scope of our approach to cohomology in this context, we will use two main tools : 1) the exactness of the inflation-restriction sequence for $n=1$, which can be proved by using elementary computations on the cochains 2) the existence of long exact sequences of cohomology groups induced by short exact sequences of $G$-modules. In particular, the Hoshchild-Serre spectral sequence will play no role in our approach.
We formulate our results for groups, but we could have opted for the point of view of Lie algebras (see the Appendix).
The main results we prove are the following :
@@
  • In a finite-dimensional theory, let $G$ be a definable connected group, and let $C$ be a definable connected Cartan subgroup of a definable connected normal subgroup $H$. Suppose that $G/N_{G}(C)$ is abelian. Then
    $G=HN_{G}(C).$
  • In a finite-dimensional theory, let $T$ be a definable connected abelian group with trivial $p$-torsion and $A$ be a definable connected $T$-module $p$-elementary as an abelian group. Suppose that $A^{T}=0$. Then $A=A_{1}\oplus\dots\oplus A_{r}$, where each $A_{i}$ is a definable connected $T$-irreducible sub-module of $A$.
In the second section, we recall some basic results about the cohomology of groups. In the third section, we consider the case of a finite nilpotent group $G$ and a finite $G$-module and we give an elementary proof of the vanishing of the first cohomology group when the submodule of invariants is trivial. In the fourth section, we generalize the cohomological results for finite groups to the case of nilpotent groups definable in a finite-dimensional theory. In the fifth section, based on the established cohomological results, we derive a weak form of the Frattini argument for connected definable Cartan subgroups. In the sixth section, we prove a definable Maschke’s theorem for definable abelian groups without $p$-torsion acting on definable abelian $p$-elementary groups. Finally, in an appendix, we explain what happens to our results when we consider Lie algebras instead of groups.
We thank Adrien Deloro for his many comments and suggestions.
2 Preliminaries on the cohomology of groups
We recall some basic facts regarding the cohomology of groups; for the sake of completeness, we give proofs when it is needed. For general references on this topic, we refer to [7] and to [10].
@@ 2 Preliminaries on the cohomology of groups
  • $C^{0}(G,A)=A$
  • For every integer $n\geq 1$, $C^{n}(G,A)$ denotes the abelian group of functions between $G^{n}$ and AA
@@
The group $G$ acts on the cochain groups of the form $C^{n}(H,A)$, for $H$ a normal subgroup, in a way that is compatible with the differentials : let $f\in C^{n}(H,A)$ and $x\in G$, then $(x\cdot f)(h_{1},\dots,h_{n})=x\cdot f(h_{1}^{x},\dots,h_{n}^{x})$.
@@
We observe that $G$ acts trivially on $H^{1}(G,A)$: let $f]\in H^{1}(G,A)$, where $f$ is a derivation, and let $x\in G$; then $(x\cdot f)(y)-f(y)=x\cdot f(y^{x})-f(y)=x\cdot(x^{-1}\cdot f(yx)+f(x^{-1}))-f(y)=f(yx)+f(1)-f(x)-f(y)=y\cdot f(x)+f(y)-f(x)-f(y)$ for all $y\in G$; in other words, $(x\cdot f)-f$ is an inner derivation and thus $x\cdot f]=[f]$.
We now turn to the question of the relations between exact sequences of $G$-modules and cohomology.
@@
Every short exact sequence of $G$-modules $0\rightarrow A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ induces a short exact sequence of complexes $0\rightarrow C(G,A)\overset{i^{*}}{\rightarrow}C(G,B)\overset{p^{*}}{\rightarrow}C(G,C)\rightarrow 0$, where $i^{*},p^{*}$ are defined by $i^{*}(f)=i\circ f$ and $p^{*}(f)=p\circ f$.
@@
Since $i$ and $p$ commute with the action of $G$, the formula for the differential allows us to show that $i^{*},p^{*}$ commute with the differential. Fix an integer $n$.
Let us show that $i^{*}$ is injective. Indeed, for $\phi\in C^{n}(G,A)$, $i^{*}(\phi)=0$ iff $i(\phi(\overline{x}))=0$ for all $\overline{x}\in G^{n}$ iff $\phi(\overline{x})=0$ for all $\overline{x}\in G^{n}$ by injectivity of $i$.
Now, we turn to the equality $\operatorname{im}(i^{*})=\ker(p^{*})$. Let $\phi\in\operatorname{im}(i^{*})\subseteq C^{n}(G,B)$; then there exists $\psi\in C^{n}(G,A)$, such that $\phi=i\circ\psi$ and hence $p^{*}(\phi)=p\circ i\circ\psi$. Consequently, for every $\overline{x}\in G^{n}$, $p(i(\psi(\overline{x}))=0$ since $\operatorname{im}(i)\subseteq\ker(p)$, and thus $\phi\in\ker(p^{*})$.
Conversely, let $\phi\in\ker(p^{*})$; then $p(\phi(\overline{x}))=0$ for every $\overline{x}\in G^{n}$. Thus $\phi(\overline{x})\in\operatorname{im}(i)$, since $\operatorname{im}(i)=\ker(p)$. Therefore for every $\overline{x}$ there exists an unique $y_{\overline{x}}\in A$, such that $i(y_{\overline{x}})=\phi(\overline{x})$ by injectivity of $i$. We can define an element of $C^{n}(G,A)$ by $\overline{x}\mapsto y_{\overline{x}}$.
Now, we prove the surjectivity of $p^{*}$. Let $s:C\rightarrow B$ be a set-theoretic section, i.e., $p\circ s=Id_{C}$. Let $\phi\in C^{n}(G,C)$; for $\overline{x}\in G$, consider $s(\phi(\overline{x}))$. Therefore, it suffices to define a cochain $\psi\in C^{n}(G,B)$ by $\psi(\overline{x})=s(\phi(\overline{x}))$ to get a map such that $p^{*}(\psi)=\phi$. ∎
Recall now an important property of the sequence of cohomology groups of a complex: to a short exact sequence of complexes, one can canonically associate a long exact sequence of cohomology groups.
Fact 2.7.
[14, Theorem 1.3.1], [9, Proposition 6.9]
Let $0\rightarrow C^{\prime}\rightarrow C\rightarrow C^{\prime\prime}\rightarrow 0$ be a short exact sequence of cochain complexes. Then there exists a long exact sequence of cohomology groups of the following form :
$\dots\rightarrow H^{n}(C^{\prime})\rightarrow H^{n}(C)\rightarrow H^{n}(C^{\prime\prime})$
$\overset{\delta_{n}}{\rightarrow}H^{n+1}(C^{\prime})\rightarrow H^{n+1}(C)\rightarrow\dots$
The morphisms $(\delta_{n})_{n\in\mathbb{Z}}$ are called connecting morphisms.
As a consequence, an exact sequence of $G$-modules induces a long exact sequence of cohomology groups.
@@
Since the inflation and restriction morphisms commute with the differentials, we obtain a sequence of complexes. In particular, we deduce a sequence $H^{1}(G/H,A^{H})\overset{\inf}{\rightarrow}H^{1}(G,A)\overset{\operatorname{res}}{\rightarrow}H^{1}(H,A)$. In fact, this sequence is exact, and it can be proved by an elementary algebraic argument.
Fact 2.9.
@@
Since $A^{H}=A$, by fact 2.9 and fact 2.10, the sequence $0\rightarrow H^{1}(G/H,A)\overset{\inf}{\rightarrow}H^{1}(G,A)\overset{\operatorname{res}}{\rightarrow}H^{1}(H,A)^{G/H}$ is exact. Therefore, it suffices to prove that
@@
3 The case of finite groups
@@
    Now, let $f:G\rightarrow A$ be a non-trivial $G$-derivation. Let $y\in G-\{1\}$, such that $f(y)\neq 0$; then $(\rho(y)-Id)$ is invertible. Let $a=(\rho(y)-Id)^{-1}\cdot f(y)$. For $x\in G$, by commutativity of $G$ and the definition of a derivation, we have $(\rho(y)-Id)\cdot f(x)=\rho(x)\cdot f(y)-f(y)$, thus, $f(x)=(\rho(y)-Id)^{-1}\cdot\rho(x)\cdot f(y)-(\rho(y)-Id)^{-1}\cdot f(y)=\rho(x)\cdot a-a$ by commutativity again. This shows that $H^{1}(G,A)=0$.
  • $\lg(A)>1$. Then, there exists a proper non-trivial $G$-module $B$, such that $\lg(B)<\lg(A)$ and $\lg(A/B)<\lg(A)$. Note that by hypothesis $B^{G}=0$. By induction on the length, we may therefore assume that $H^{1}(G,B)=0$. The long exact cohomology sequence will now come into play. Indeed, we have a short exact sequence of $G$-modules $0\rightarrow B\rightarrow A\rightarrow A/B\rightarrow 0$ which, by virtue of Fact 2.7, induces the following long exact cohomology sequence :
@@
    $\overset{\delta_{0}}{\rightarrow}H^{1}(G,B)\rightarrow H^{1}(G,A)\rightarrow H^{1}(G,A/B)\rightarrow\dots$
    By hypothesis, $A^{G}=B^{G}=0$ and by induction hypothesis, $H^{1}(G,B)=0$, so the exactness of the sequence then gives us $(A/B)^{G}=0$; we may therefore apply once more the induction hypothesis and use the exactness of the long cohomology sequence to establish $H^{1}(G,A)=0$.
@@
  • $A^{Z(G)}=A$. Lemma 2.11 gives $H^{1}(Z(G),A)^{G/Z(G)}=0$; so, the inflation-restriction exact sequence entails $H^{1}(G/Z(G),A)\simeq H^{1}(G,A)$ and we are done by induction.
@@
To extend the proof to the category of definable $G$-modules in a first order theory, we need two main ingredients : 1) a notion of irreducible module that allows inductive reasoning on the lenght 2) an appropriate version of Schur’s Lemma regarding this irreducible category. In the next section, we will see that finite-dimensional theories satisfy these requirements.
4 Finite-dimensional nilpotent groups
@@
A strongly minimal theory, the theory of a group of finite Morley rank and of an o-minimal structure are examples of finite-dimensional theories (for o-minimal structures, we must restrict to the collection of definable sets and interpretable group quotients). This model-theoretic framework is suitable for inductive reasoning when working with groups.
Fact 4.2.
@@
  • First, suppose $\lg(A)=1$, i.e., $A$ is irreducible. We denote by $\rho:G\rightarrow\operatorname{Aut}_{def}(A)\subseteq\operatorname{End}_{def}(A)$ the action. By Fact 4.3, we can linearize : there exists a definable skew-field $\mathbb{K}$ such that $\rho(G)\leq\mathbb{K}$. Up to considering $Z(\mathbb{K})$, which is also definable infinite, we may assume that $\mathbb{K}$ is commutative. Now, the proof goes as in the Lemma 3.2.
@@
A strongly minimal theory, the theory of an o-minimal structure, the theory of a group of finite Morley rank belong to this subclass. In a finite-dimensional theory, a definable irreducible $G$-module has no proper infinite definable submodule. Moreover, an infinite definable nilpotent group has infinite center (see [1, Lemma 5.1, I]).
@@
  • $A^{Z(G)^{\circ}}$ is finite. In fact, $A^{Z(G)^{\circ}}=0$ since $A^{G}=0$ and $G$ is connected. Notice that $H^{1}(G/Z(G)^{\circ},A^{Z(G)^{\circ}})=0=H^{1}(Z(G)^{\circ},A)$ by Lemma 4.4 and we are done by fact 2.9.
@@
$\overset{\delta_{0}}{\rightarrow}H^{1}(G,V)\dots$
@@
5 A weak Frattini argument for definable connected Cartan subgroups
In the theory of finite groups, $p$-Sylow subgroups play a fundamental role and often allow to obtain structural results about the ambient group. In particular, the Frattini argument states that, for a finite group $G$, a normal subgroup $H$, and a $p$-Sylow subgroup $S$ of $H$, we have the following :
$G=HN_{G}(S),$
where $N_{G}(S)$ denotes the normalizer of $S$ in $G$. In fact, the image of $S$ under an inner automorphism $inn(g)$ remains a $p$-Sylow subgroup of $H$; by the conjugacy of $p$-Sylow subgroups in $H$, there exists $h\in H$, such that $S^{h}=S^{g}$, that is, such that $h^{-1}g\in N_{G}(S)$.
Outside the world of finite groups, Cartan subgroups, i.e., nilpotent self-normalizing subgroups can be seen as a good proxy for $p$-Sylow subgroups. In general, the existence and the conjugation of Cartan subgroups are not guaranteed. Nevertheless, in a finite-dimensional theory, if there exist definable connected Cartan subgroups, the conclusion of the Frattini argument may still be true without any conjugation assumption. We follow the cohomological treatment for finite-dimensional Lie algebras given by Barnes in [4].
Definition 5.1.
In a finite-dimensional theory, let $G$ be a definable connected group and let $C$ be a definable connected nilpotent subgroup such that
$N_{G}(C)=C.$
We then say that $C$ is a definable connected Cartan subgroup.
Proposition 5.2.
In a finite-dimensional theory, let $G$ be a definable connected group, and let $C$ be a definable connected Cartan subgroup of a definable connected normal subgroup $H$. Suppose that $G/N_{G}(C)$ is abelian. Then
$G=HN_{G}(C).$
Proof.
We set $N=N_{G}(C)$ and we consider the definable connected $C$-module $A=G/N$.
Let us prove that $A^{C}=0$. Indeed, let $\overline{a}\in A$ be such that $c\cdot\overline{a}=\overline{a}$ for all $c\in C$; then $a^{-1}a^{c}\in H\cap N$. But $H$ is normal subgroup and $N_{H}(C)=C$, so $a^{-1}a^{c}\in C$ for all $c$, hence $a\in N_{G}(C)$.
Now, consider the $C$-module
$(G/N)/((HN)/N)\simeq G/(HN).$
If this module were nontrivial, Corollary 4.7 would imply that $C$ acts non-trivially, which contradicts the fact that $H$ contains $C$. Therefore, we conclude that
$G=HN.\qed$

### v2

# On the cohomology of finite-dimensional nilpotent groups and Lie rings [^1] Keywords: nilpotent group, nilpotent Lie algebra, cohomology, finite-dimensional theory MSC 2020 : 17B56, 03C60

@@ Abstract
We give some vanishing results for the first classical cohomology group in the case of finite-dimensional nilpotent groups and Lie rings. Based on the established cohomological results, we derive some structural consequences : we prove a form of Frattini’s argument for definable connected Cartan subrings and we give a definable version of Maschke’s theorem for the action of a definable connected $p$-divisible abelian group.
@@ 1 Introduction
Under the impetus of Hochschild and Serre (see [8] and [9]), techniques from homological algebra have been successfully applied to the study of modules in the category of groups and of Lie algebras. For (finite) groups, cohomological methods play a prominent role in the classification of group extensions and had striking applications in number theory through Galois cohomology (see [10]). In the case of Lie algebras, a number of structural results for finite-dimensional (in the usual linear sense) solvable Lie algebras were obtained by Barnes by using cohomological tools (see [2] and [3]). The approach of Barnes is particulary relevant in our case; more precisely, in [2] he proved that the cohomology groups associated to a nilpotent Lie algebra of finite linear dimension vanishe if the submodule of invariants is trivial. The same result is true when one replaces nilpotent Lie algebras by nilpotent finite groups or algebraic groups, nevertheless, a general and uniform treatement for this kind of vanishing results seems to be lacking.
Model-theory could provide this kind of general framework and we generalize Barnes’ methods to the context of groups and Lie rings (Lie algebras devoided of their vector space structure) equipped with some kind of dimension function on the definable subsets, which generalizes the properties of the linear dimension function and cardinality. From a model-theoretic point of view, we will consider groups and Lie rings that are definable in a finite-dimensional theory (see below for a precise definition). Finite groups, algebraic groups over an algebraically closed field, semi-algebraic groups over the reals are some striking examples of such "finite-dimensional" groups. On the other hand, finite-dimensional Lie algebras are the main instances of "finite-dimensional" Lie rings. Nevertheless, we decided to focus on the abstract setting of Lie rings since this kind of structures have recently captured the attention of the model-theoretic community (see [5] and [6] for example) and we hope what is presented here will contribute to this line of research. As an additional upshot, we revisite some standard algebraic results such as Frattini-like arguments and Maschke-like theorems.
Before stating our results, we say some words regarding the scope of our approach to cohomology. We will use two main tools : 1) the exactness of the inflation-restriction sequence for $n=1$, which can be proved by using elementary computations on the cochains 2) the existence of "long" exact sequences of cohomology groups induced by short exact sequences of $G$-modules. In particular, the Hoshchild-Serre spectral sequence will play no role in our approach.
Regarding cohomology, we prove the following :
Theorem 1.1.
  1. 1.
@@
We formulate our results for groups but we could have opted for the point of view of Lie rings. On the basis of our cohomological results, we give the two following algebraic applications:
Theorem 1.2.
  1. 1.
    (definable Maschke’s theorem) In a finite-dimensional theory, let $T$ be a definable connected abelian group with trivial $p$-torsion and $A$ be a definable connected $T$-module $p$-elementary as an abelian group. Suppose that $A^{T}=0$. Then $A=A_{1}\oplus\dots\oplus A_{r}$, where each $A_{i}$ is a definable connected $T$-irreducible sub-module of $A$.
  2. 2.
    (Frattini’s argument for Lie rings) In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected Lie ring, and let $\mathfrak{c}$ be a definable connected Cartan subring of a definable connected ideal $\mathfrak{i}$.Then
    $\mathfrak{g}=\mathfrak{i}+N_{\mathfrak{g}}(\mathfrak{c}).$
In the second section, we recall some basic results about the cohomology of groups and after reviewing the case of finite nilpotent groups, we give an elementary proof of the vanishing of the first cohomology group when the submodule of invariants is trivial for a definable connected nilpotent group in a finite-dimensional theory. In the third section, we give our first algebraic application and we prove a definable version of Maschke’s theorem for the action of a definable connected $p$-divisible abelian group. In the fourth section, we pass to the case of Lie rings and in particular, we give a version of Frattini’s argument for definable connected Cartan subrings.
We thank Adrien Deloro for his many comments and suggestions. We are also grateful to Moreno Invitti for his key observation regarding long exact cohomological sequences.
2 Cohomological results
2.1 Preliminaries on the cohomology of groups
We recall some basic facts regarding the cohomology of groups. Although the results presented in this subsection are fairly standard, we decided to provide complete elementary algebraic proofs that could easily generalize to the case of abstract Lie rings for which general cohomological results may fail. For general references on this topic, we refer to [8] and to [10].
@@ 2.1 Preliminaries on the cohomology of groups
  • $C^{0}(G,A)=A$.
  • For every integer $n\geq 1$, $C^{n}(G,A)$ denotes the abelian group of functions between $G^{n}$ and $A$.
@@
The group $G$ acts on the cochain groups of the form $C^{n}(H,A)$, for $H$ a normal subgroup, in a way that is compatible with the differentials :
let $f\in C^{n}(H,A)$ and $x\in G$, then $(x\cdot f)(h_{1},\dots,h_{n})=x\cdot f(h_{1}^{x},\dots,h_{n}^{x})$.
@@
We observe that $G$ acts trivially on $H^{1}(G,A)$: let $f]\in H^{1}(G,A)$, where $f$ is a derivation, and let $x\in G$; then
$(x\cdot f)(y)-f(y)=x\cdot f(y^{x})-f(y)$
$=x\cdot(x^{-1}\cdot f(yx)+f(x^{-1}))-f(y)$
$=f(yx)+f(1)-f(x)-f(y)=y\cdot f(x)+f(y)-f(x)-f(y)$
$=y\cdot f(x)-f(x)$
for all $y\in G$; in other words, $(x\cdot f)-f$ is an inner derivation and thus $x\cdot f]=[f]$.
We now turn to the question of the relations between exact sequences of $G$-modules and long exact sequence of cohomology groups.
@@
Every short exact sequence of $G$-modules $0\rightarrow A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ induces the following short exact sequence :
$0\rightarrow C^{1}(G,A)\overset{i^{*}}{\rightarrow}C^{1}(G,B)\overset{p^{*}}{\rightarrow}C^{1}(G,C),$
where $i^{*},p^{*}$ are defined by $i^{*}(f)=i\circ f$ and $p^{*}(f)=p\circ f$. Moreover, $i^{*}$ and $p^{*}$ commute with the differential.
@@
Since $i$ and $p$ commute with the action of $G$, the formula for the differential allows us to show that $i^{*},p^{*}$ commute with it.
Let us show that $i^{*}$ is injective. Indeed, for $\phi\in C^{1}(G,A)$, $i^{*}(\phi)=0$ iff $i(\phi(x))=0$ for all $x\in G$ iff $\phi(x)=0$ for all $x\in G$ by injectivity of $i$.
Now, we turn to the equality $\operatorname{im}(i^{*})=\ker(p^{*})$. Let $\phi\in\operatorname{im}(i^{*})\subseteq C^{1}(G,B)$; then there exists $\psi\in C^{1}(G,A)$, such that $\phi=i\circ\psi$ and hence $p^{*}(\phi)=p\circ i\circ\psi$. Consequently, for every $x\in G$, $p(i(\psi(x))=0$ since $\operatorname{im}(i)\subseteq\ker(p)$, and thus $\phi\in\ker(p^{*})$.
Conversely, let $\phi\in\ker(p^{*})$; then $p(\phi(x))=0$ for every $x\in G$. Thus $\phi(x)\in\operatorname{im}(i)$, since $\operatorname{im}(i)=\ker(p)$. Therefore for every $x$ there exists an unique $y_{x}\in A$, such that $i(y_{x})=\phi(x)$ by injectivity of $i$. Now, it suffices to consider the following element of $C^{1}(G,A)$ : $\psi:x\mapsto y_{x}$. ∎
Corollary 2.7.
Every short exact sequence of $G$-modules $0\rightarrow A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ induces the following short exact sequence :
$0\rightarrow Der(G,A)\overset{i^{*}}{\rightarrow}Der(G,B)\overset{p^{*}}{\rightarrow}Der(G,C),$
where $i^{*},p^{*}$ are defined by $i^{*}(f)=i\circ f$ and $p^{*}(f)=p\circ f$.
Proof.
Since $i$ and $p$ are $G$-morphisms, the sequence is well-defined. Moreover, let $\phi\in\ker(p^{*})$ and consider the application $\psi:G\rightarrow A,x\mapsto y_{x}$, where $i(y_{x})=\phi(x)$. Then, for all $x,y\in G$, we have $i(x\cdot\psi(y)+\psi(x))=x\cdot i(\psi(y))+i(\psi(x))=x\cdot\phi(y)+\phi(x)=\phi(xy)=i(\psi(xy))$, so, by injectivity of $i$, $\psi(xy)=x\cdot\psi(y)+\psi(x)$. ∎
One important feature of cohomology is the possibility of getting a "long" exact sequence of cohomology groups from a short exact sequence of $G$-modules.
Fact 2.8.
Let $0\rightarrow A_{0}\rightarrow A_{1}\rightarrow A_{3}\rightarrow 0$ be a short exact sequence of $G$-modules. Then there exists a "long" exact sequence of cohomology groups of the following form :
$A_{0}^{G}\rightarrow A_{1}^{G}\rightarrow A_{2}^{G}$
$\overset{\delta_{0}}{\rightarrow}H^{1}(G,A_{0})\rightarrow H^{1}(G,A_{1})\rightarrow\ H^{1}(G,A_{2})$
The morphism $(\delta_{0})$ is called a connecting morphism.
Proof.
It suffices to apply the Snake Lemma to the following commutative diagram, where $\phi_{i}:A_{i}\rightarrow Der(G,A_{i}),a_{i}\mapsto(g\mapsto g\cdot a_{i}-a_{i}),i=0,1,2$ :
[…]
@@
Since the inflation and restriction morphisms commute with the differentials, we can pass to the cohomology groups and we get a sequence $H^{1}(G/H,A^{H})\overset{\inf}{\rightarrow}H^{1}(G,A)\overset{\operatorname{res}}{\rightarrow}H^{1}(H,A)$. In fact, this sequence is exact, and it can be proved by an elementary algebraic argument.
Fact 2.10.
@@
Since $A^{H}=A$, by Fact 2.10 and Fact 2.11, the sequence $0\rightarrow H^{1}(G/H,A)\overset{\inf}{\rightarrow}H^{1}(G,A)\overset{\operatorname{res}}{\rightarrow}H^{1}(H,A)^{G/H}$ is exact. Therefore, it suffices to prove that
@@
2.2 The case of finite groups
@@
    Now, let $f:G\rightarrow A$ be a non-trivial $G$-derivation. Let $y\in G-\{1\}$, such that $f(y)\neq 0$; then $(\rho(y)-Id)$ is invertible. Let $a=(\rho(y)-Id)^{-1}\cdot f(y)$. For $x\in G$, by commutativity of $G$ and the definition of a derivation, we have $(\rho(y)-Id)\cdot f(x)=\rho(x)\cdot f(y)-f(y)$, thus,
    $f(x)=(\rho(y)-Id)^{-1}\cdot\rho(x)\cdot f(y)-(\rho(y)-Id)^{-1}\cdot f(y)=\rho(x)\cdot a-a$
    by commutativity again. This shows that $H^{1}(G,A)=0$.
  • $\lg(A)>1$. Then, there exists a proper non-trivial $G$-module $B$, such that $\lg(B)<\lg(A)$ and $\lg(A/B)<\lg(A)$. Note that by hypothesis $B^{G}=0$. By induction on the length, we may therefore assume that $H^{1}(G,B)=0$. The long exact cohomology sequence will now come into play. Indeed, we have a short exact sequence of $G$-modules $0\rightarrow B\rightarrow A\rightarrow A/B\rightarrow 0$ which, by virtue of Fact 2.8, induces the following long exact cohomology sequence :
@@
    $\overset{\delta_{0}}{\rightarrow}H^{1}(G,B)\rightarrow H^{1}(G,A)\rightarrow H^{1}(G,A/B)$
    By hypothesis, $A^{G}=B^{G}=0$ and by induction hypothesis, $H^{1}(G,B)=0$, so the exactness of the sequence then gives us that $(A/B)^{G}=0$; we may therefore apply once more the induction hypothesis and use the exactness of the long cohomology sequence to establish that $H^{1}(G,A)=0$.
@@
  • $A^{Z(G)}=A$. Then Lemma 2.12 gives $H^{1}(Z(G),A)^{G/Z(G)}=0$; so, the inflation-restriction exact sequence entails $H^{1}(G/Z(G),A)\simeq H^{1}(G,A)$ and we are done by induction.
@@
To extend the proof to the category of definable $G$-modules in a first order theory, we need two main ingredients : 1) a notion of irreducible module that allows inductive reasoning on the lenght 2) an appropriate version of Schur’s Lemma regarding this irreducible category. In the next subsection, we will see that finite-dimensional theories satisfy these requirements.
2.3 Finite-dimensional nilpotent groups
@@
  • (Low fibration) If $f:X\longrightarrow Y$ is an interpretable map, such that $\dim(f^{-1}(y))\leq d$ for all $y\in Y$, then $\dim(X)\leq\dim(Y)+d$.
@@
The theory of a group of finite Morley rank and of an o-minimal structure are examples of finite-dimensional theories (for o-minimal structures, we must restrict to the collection of definable sets and interpretable group quotients). This model-theoretic framework is suitable for inductive reasoning when working with groups.
Fact 2.18.
@@
  • First, suppose $\lg(A)=1$, i.e., $A$ is irreducible. We denote by $\rho:G\rightarrow\operatorname{Aut}_{def}(A)\subseteq\operatorname{End}_{def}(A)$ the action. By Fact 2.19, we can linearize : there exists a definable skew-field $\mathbb{K}$ such that $\rho(G)\leq\mathbb{K}$. Up to considering $Z(\mathbb{K})$, which is also definable infinite, we may assume that $\mathbb{K}$ is commutative. Now, the proof goes as in the Lemma 2.14.
@@
The theory of an o-minimal structure and the theory of a group of finite Morley rank belong to this subclass. In a finite-dimensional theory, a definable irreducible $G$-module has no proper infinite definable submodule. Moreover, an infinite definable nilpotent group has infinite center (see [1, Lemma 5.1, I]).
@@
  • $A^{Z(G)^{\circ}}$ is finite. In fact, $A^{Z(G)^{\circ}}=0$ since $A^{G}=0$ and $G$ is connected. Notice that $H^{1}(G/Z(G)^{\circ},A^{Z(G)^{\circ}})=0=H^{1}(Z(G)^{\circ},A)$ by Lemma 2.20 and we are done by Fact 2.10.
@@
$\overset{\delta_{0}}{\rightarrow}H^{1}(G,V)$